# GeoClip: Geometry-Aware Clipping for Differentially Private SGD

### v2

@@ 2 General Framework
The above result generalizes Theorem 2 in Pichapati et al. [9]; proof details are in the supplement.
@@
We now show that the remaining contribution to the clipping error term $\operatorname{Tr}\left(M_{t}^{\top}M_{t}\Sigma_{t}\right)$ in fact serves as an upper bound on the probability that the gradient is clipped (i.e., $|\omega|>1$). We do so using Markov’s inequality as detailed below:
@@
The proof of Theorem˜2 is in the supplement.
@@ 3 Algorithm Overview
1:Eigenvectors $V\in\mathbb{R}^{d\times k}$, eigenvalues $\Lambda\in\mathbb{R}^{k\times k}$, gradient $\tilde{g}$, mean $a$, factor $\beta_{3}$, rank $k$
2:Center: $z\leftarrow\tilde{g}-a$
3:Form augmented matrix: $U_{\text{aug}}\leftarrow[V\ \ z]$
@@
5:Perform SVD: $Z=USR^{\top}$
6:Set $V_{\text{new}}\leftarrow$ first $k$ columns of $U$
7:Set $\Lambda_{\text{new}}\leftarrow$ squares of the first $k$ singular values in $S$
8:Return: $V_{\text{new}},\Lambda_{\text{new}}$
Low-Rank PCA. When the dimensionality is high, computing and storing the full gradient covariance matrix becomes impractical. To address this, we propose a method to maintain a low-rank approximation using a simple and efficient procedure we refer to as Streaming Rank-$k$ PCA (Algorithm 2). Specifically, we maintain an approximate eigendecomposition of the covariance in the form $U\Lambda U^{\top}$, where $U\in\mathbb{R}^{d\times k}$ contains the top-$k$ eigenvectors and $\Lambda\in\mathbb{R}^{k\times k}$ the corresponding eigenvalues, where $k\ll d$. Upon receiving a new gradient $\tilde{g}$, we center it using the running mean $a$, yielding $z=\tilde{g}-a$, and perform a weighted update to the covariance:
$\Sigma_{\text{new}}$ $=\beta_{3}U\Lambda U^{\top}+(1-\beta_{3})zz^{\top}=[U\;\;z]\begin{bmatrix}\beta_{3}\Lambda&0\\ 0&1-\beta_{3}\end{bmatrix}[U\;\;z]^{\top}$ (21)
@@
$Z=[U\;\;z]\cdot\mathrm{diag}(\sqrt{\beta_{3}\lambda_{1}},\dots,\sqrt{\beta_{3}\lambda_{k}},\sqrt{1-\beta_{3}})\in\mathbb{R}^{d\times(k+1)}$ (22)
We then perform an SVD on $Z$ and retain the top $k$ singular vectors and squared singular values as the updated eigenvectors and eigenvalues. We note that this computation takes $\mathcal{O}(dk^{2}+k^{3})$ time [14], highlighting that it is only linear in $d$. The rest of the procedure follows Algorithm 1 by replacing the moving average over the full covariance matrix in line 12 of Algorithm˜1 with the low-rank approximation described above. Another necessary change to the algorithm is the following: since $U_{t}$ is no longer a square matrix, the transformation $M_{t}$ takes the gradient into the lower $k$-dimensional space to clip and add noise, such that the resulting $M_{t}^{\text{inv}}$ returns to the full $d$-dimensional space. The complete version of this variant is provided as Algorithm 3 in the supplementary material.
@@ 4 Experimental Results
We present empirical results demonstrating the benefits of our proposed GeoClip framework compared to AdaClip, quantile-based clipping, and standard DP-SGD. For quantile-based clipping, we use the median of per-sample norms, which has been shown to perform well across various learning tasks [10]. All $(\varepsilon,\delta)$-DP guarantees are computed using the Connect-the-Dots accountant [15]. All experiments were conducted on Google Colab using CPU resources.
@@ 4.1 Synthetic Dataset
GeoClip is designed to improve convergence, particularly in the presence of feature correlation. To empirically demonstrate this, we evaluate it on a synthetic Gaussian dataset with 20,000 samples and 10 features—five of which are correlated, while the remaining five are independent. We train a linear regression model using various private training methods for 10 epochs with a batch size of 1024, tuning the learning rate for each method to ensure stable convergence. As shown in Figure 1 (left), GeoClip converges as early as epoch 2, while the next best method—quantile-based clipping—requires nearly twice as many epochs. Figure 1 (right) plots the privacy cost ($\varepsilon$) versus epoch, showing how faster convergence helps minimize overall privacy cost.
@@ 4.4 Low-Rank PCA Results
To evaluate our method on low-rank PCA algorithm (Algorithm 2), we construct a synthetic binary classification dataset with 20,000 samples and 400 Gaussian features, where 50 are correlated and 350 are uncorrelated. Labels are generated by applying a linear function to the features, adding Gaussian noise, and thresholding the sigmoid output. We train a logistic regression model with 802 trainable parameters. We then apply our method using a low-rank PCA approximation with rank 50. As shown in the left panel of Figure 2, our approach converges faster than competing methods even with this low-rank approximation.
We also evaluate our method on the USPS dataset [21] using logistic regression with 2,570 trainable parameters (256 input features $\times$ 10 classes + 10 biases). The USPS dataset contains 9,298 grayscale handwritten digit images (0–9), each of size $16\times 16$ pixels. This compact benchmark is commonly used for evaluating digit classification models. For this dataset, we apply a low-rank PCA approximation with rank 100. Results are shown in the right panel of Figure 2. As with the synthetic dataset, GeoClip with low-rank PCA also achieves faster convergence on USPS compared to baseline methods. We provide additional $\varepsilon$-vs-iteration plots for both datasets, illustrating how faster convergence reduces overall privacy cost in the supplementary material.
@@ Appendix C GeoClip with Low-Rank PCA
Computing and storing the full gradient covariance matrix becomes infeasible in high dimensions. To address this, we propose a low-rank approximation method in Algorithm 3, which incorporates a rank-$k$ PCA step described in the Streaming Rank-$k$ PCA algorithm. This procedure efficiently maintains an approximate eigendecomposition $U_{t}\Lambda_{t}U_{t}^{\top}$, where $U_{t}\in\mathbb{R}^{d\times k}$ contains the top-$k$ eigenvectors and $\Lambda_{t}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{k})\in\mathbb{R}^{k\times k}$ holds the corresponding eigenvalues at iteration $t$, with $k\ll d$.
Given a new gradient $\tilde{g}_{t}$, we center it using the running mean $a_{t+1}$ to obtain $z=\tilde{g}_{t}-a_{t+1}$, and update the covariance as:
$\Sigma_{\text{new}}=\beta_{3}U_{t}\Lambda_{t}U_{t}^{\top}+(1-\beta_{3})zz^{\top}=[U_{t}\;\;z]\begin{bmatrix}\beta_{3}\Lambda_{t}&0\\ 0&1-\beta_{3}\end{bmatrix}[U_{t}\;\;z]^{\top}.$ (77)
Rather than forming this matrix directly, we compute its square root:
$Z=[U_{t}\;\;z]\cdot\mathrm{diag}(\sqrt{\beta_{3}\lambda_{1}},\dots,\sqrt{\beta_{3}\lambda_{k}},\sqrt{1-\beta_{3}}),$ (78)
and perform an SVD on $Z$, retaining the top $k$ components to obtain $U_{t+1}$ and $\Lambda_{t+1}$.
Since $U_{t}$ is not square, $M_{t}\in\mathbb{R}^{k\times d}$ projects gradients into a $k$-dimensional subspace for clipping and noise addition, while $M_{t}^{\text{inv}}\in\mathbb{R}^{d\times k}$ maps them back to $\mathbb{R}^{d}$.
@@
Appendix D Privacy Cost vs. Iteration (Corresponding to Figure 2)
Figure 4 shows the privacy cost ($\varepsilon$) versus iteration curves associated with the accuracy-vs-iteration plots in Figure 2 (Section 4.4, Low-Rank PCA Results), which is repeated in this PDF as Figure 3 for easier reference. These plots highlight how faster convergence reduces overall privacy cost. For example, on the USPS dataset (Figure 3, right), GeoClip achieves high accuracy within the first few iterations, requiring only $\varepsilon\approx 0.15$, whereas quantile-based clipping takes about 24 iterations to reach similar accuracy, incurring a higher privacy cost of $\varepsilon\approx 0.5$.
Figure 4: The left plot shows the overall privacy budget $\varepsilon$ spent on training the synthetic Gaussian dataset with 400 features for $\delta=10^{-5}$, while the right plot shows the same for the USPS dataset with $\delta=10^{-5}$. These plots apply to all four algorithms, which are tuned to achieve the same privacy level for a given number of iterations.

### v3

@@ 2 General Framework
The above result generalizes Theorem 2 in Pichapati et al. [9]; proof details are in Appendix A.
@@
We now show that the remaining contribution to the clipping error term $\operatorname{Tr}\left(M_{t}^{\top}M_{t}\Sigma_{t}\right)$ in fact serves as an upper bound on the probability that the gradient is clipped (i.e., $\|\omega\|>1$). We do so using Markov’s inequality as detailed below:
@@
The proof of Theorem˜2 is provided in Appendix B.
@@ 3 Algorithm Overview
2:Center: $z\leftarrow\tilde{g}_{t}-a_{t+1}$
3:Form augmented matrix: $U_{\text{aug}}\leftarrow[U_{t}\ \ z]$
@@
5:Perform SVD: $Z=VSR^{\top}$
6:Set $U_{t+1}\leftarrow$ first $k$ columns of $V$
8:Return: $U_{t+1},\Lambda_{t+1}$
Low-Rank PCA. When the dimensionality is high, computing and storing the full gradient covariance matrix becomes impractical. To address this, we propose a method to maintain a low-rank approximation using a simple and efficient procedure we refer to as Streaming Rank-$k$ PCA (Algorithm 2). Specifically, we maintain an approximate eigendecomposition of the covariance in the form $U_{t}\Lambda_{t}U_{t}^{\top}$, where $U_{t}\in\mathbb{R}^{d\times k}$ contains the top-$k$ eigenvectors and $\Lambda_{t}\in\mathbb{R}^{k\times k}$ the corresponding eigenvalues, where $k\ll d$. Upon receiving a new gradient $\tilde{g}_{t}$, we center it using the running mean $a_{t+1}$, yielding $z=\tilde{g}_{t}-a_{t+1}$, and perform a weighted update to the covariance:
$\Sigma_{t+1}$ $=\beta_{3}U_{t}\Lambda_{t}U_{t}^{\top}+(1-\beta_{3})zz^{\top}=[U_{t}\;\;z]\begin{bmatrix}\beta_{3}\Lambda_{t}&0\\ 0&1-\beta_{3}\end{bmatrix}[U_{t}\;\;z]^{\top}$ (21)
@@
$Z=[U_{t}\;\;z]\cdot\mathrm{diag}(\sqrt{\beta_{3}\lambda_{1}},\dots,\sqrt{\beta_{3}\lambda_{k}},\sqrt{1-\beta_{3}})\in\mathbb{R}^{d\times(k+1)}.$ (22)
We then perform an SVD on $Z$ and retain the top $k$ singular vectors and squared singular values as the updated eigenvectors and eigenvalues. We note that this computation takes $\mathcal{O}(dk^{2}+k^{3})$ time [14], highlighting that it is only linear in $d$. The rest of the procedure follows Algorithm 1 by replacing the moving average over the full covariance matrix in line 12 of Algorithm˜1 with the low-rank approximation described above. Another necessary change to the algorithm is the following: since $U_{t}$ is no longer a square matrix, the transformation $M_{t}$ takes the gradient into the lower $k$-dimensional space to clip and add noise, such that the resulting $M_{t}^{\text{inv}}$ returns to the full $d$-dimensional space. The complete version of this variant is provided as Algorithm 3 in Appendix C.
@@ 4 Experimental Results
We present empirical results demonstrating the benefits of our proposed GeoClip framework compared to AdaClip, quantile-based clipping, and standard DP-SGD. For quantile-based clipping, we use the median of per-sample norms, which has been shown to perform well across various learning tasks [10]. All $(\varepsilon,\delta)$-DP guarantees are computed using the Connect-the-Dots accountant [15]. All experiments were conducted on Google Colab using CPU resources. The code implementation is available on GitHub [16].
@@ 4.1 Synthetic Dataset
GeoClip is designed to improve convergence, particularly in the presence of feature correlation. To empirically demonstrate this, we evaluate it on a synthetic Gaussian dataset with 20,000 samples and 10 features—five of which are correlated, while the remaining five are independent. To obtain the 5 correlated features, we first generate an $20{,}000\times 5$ matrix $Z$ and a $5\times 5$ matrix $A$, each with entries drawn independently from the standard normal distribution. The correlated features are obtained by multiplying $Z$ by $A$. The remaining 5 features are drawn independently from a standard multivariate normal distribution. The full feature matrix $X$ is constructed by concatenating these two blocks. The target $y$ is generated using a linear function with Gaussian noise: $y=Xw+b+\epsilon$, where $w\sim\mathcal{N}(0,I_{10})$ is a weight vector, $b\sim\mathcal{N}(0,1)$ is a scalar bias term, and $\epsilon\sim\mathcal{N}(0,0.01^{2})$ is i.i.d. noise. We train a linear regression model using various private training methods for 10 epochs with a batch size of 1024, tuning the learning rate for each method to ensure stable convergence. As shown in Figure 1 (left), GeoClip converges as early as epoch 2, while the next best method—quantile-based clipping—requires nearly twice as many epochs. Figure 1 (right) plots the privacy cost ($\varepsilon$) versus epoch, showing how faster convergence helps minimize overall privacy cost.
Although GeoClip introduces some computational and memory overhead, this cost is well justified by its empirical gains. GeoClip converges considerably faster than baseline methods, requiring fewer training iterations to reach comparable utility. While each iteration is slightly more expensive than in other baselines, this cost is partially offset by the reduced number of iterations. Combined with improved privacy, these benefits make the additional overhead a worthwhile trade-off.
@@ 4.4 Low-Rank PCA Results
To evaluate our low-rank PCA algorithm (Algorithm 2), we construct a synthetic binary classification dataset with 20,000 samples and 400 Gaussian features, where 50 are correlated and 350 are uncorrelated. The feature matrix is constructed in the same manner as in Section 4.1. Labels are generated by applying a linear function to the features, adding Gaussian noise, and thresholding the sigmoid output. We train a logistic regression model with 802 parameters and compare our method against existing benchmarks using a rank-50 low-rank PCA approximation. As shown in the left panel of Figure 2, our approach converges faster than competing methods, even with this low-rank approximation.
We also evaluate our method on the USPS dataset [22] using logistic regression with 2,570 trainable parameters (256 input features $\times$ 10 classes + 10 biases). The USPS dataset contains 9,298 grayscale handwritten digit images (0–9), each of size $16\times 16$ pixels. This compact benchmark is commonly used for evaluating digit classification models. For this dataset, we apply a low-rank PCA approximation with rank 100. Results are shown in the right panel of Figure 2. As with the synthetic dataset, GeoClip with low-rank PCA also achieves faster convergence on USPS compared to baseline methods. We provide $\varepsilon$-vs-iteration plots for both datasets, illustrating how faster convergence reduces overall privacy cost in Appendix D.
We compare the accuracy of GeoClip using Algorithm 1 and Algorithm 2 with $k=50$ on the same synthetic dataset, with the results shown in Figure 3 in Appendix D. The results demonstrate that even with $k=50$, Algorithm 2 achieves accuracy comparable to Algorithm 1, which computes the full covariance.
@@
We also conduct an ablation study on $\gamma$, $h_{1}$, and $h_{2}$ using the USPS dataset under the same experimental settings. The results are presented in Appendix E. The findings show that varying $\gamma$ results in only minor performance changes. Likewise, $h_{1}$ has no noticeable impact, and $h_{2}$ introduces small variations.
@@ 5 Conclusions
Limitations.
The linear transformation involved in GeoClip requires an additional computation via an eigendecomposition. Our low-rank approximation addresses that to some extent. The algorithm introduces additional hyperparameters (most notably $h_{2}$) compared to standard DP-SGD, which must be tuned for optimal performance. Our experiments have been performed on a limited collection of datasets; additional testing is needed to see how our algorithm performs in more generality.
Broader impact.
As discussed in the Introduction, our work is motivated by societal concerns, with a focus on improving the theoretical limits of differentially private optimization.
Acknowledgements
We thank the anonymous reviewers for their valuable feedback, which significantly improved the quality of this paper. This work was supported by the National Science Foundation under Grant Nos. CIF-1901243, CIF-2312666, and CIF-2007688.
@@ Appendix C GeoClip with Low-Rank PCA
Computing and storing the full gradient covariance matrix becomes infeasible in high dimensions. To address this, we propose a low-rank approximation method in Algorithm 3, which incorporates a rank-$k$ PCA step described in the Streaming Rank-$k$ PCA algorithm.
@@
Appendix D Additional Plots Related to Section 4.4
The middle and left panels of Figure 3 show the privacy cost ($\varepsilon$) versus iteration curves corresponding to the accuracy–iteration plots in Figure 2 (Section 4.4, Low-Rank PCA Results). These plots highlight how faster convergence reduces overall privacy cost. For example, on the USPS dataset (Figure 2), GeoClip achieves high accuracy within the first few iterations, requiring only $\varepsilon\approx 0.15$, whereas quantile-based clipping takes about 24 iterations to reach similar accuracy, incurring a higher privacy cost of $\varepsilon\approx 0.5$.
We compare the accuracy of GeoClip using Algorithm 1 and Algorithm 2 with $k=50$ on the same synthetic dataset described in Section 4.4, as shown in the right panel of Figure 3. The results show that even with $k=50$, Algorithm 2 achieves accuracy comparable to Algorithm 1, which computes the full covariance.
Figure 3: The left plot shows the overall privacy budget $\varepsilon$ spent on training the synthetic Gaussian dataset with 400 features for $\delta=10^{-5}$, while the middle plot shows the same for the USPS dataset with $\delta=10^{-5}$. These plots apply to all four algorithms, which are tuned to achieve the same privacy level for a given number of iterations. The right panel compares GeoClip’s performance using the full-covariance and low-rank approximation algorithms.
Appendix E Ablation Study
We conduct an ablation study on $\gamma$, $h_{1}$, and $h_{2}$ using the USPS dataset, under the same settings as in Section 4.4. As shown, varying $\gamma$ does not significantly affect performance. The results indicate no noticeable impact from $h_{1}$, and only minor variations with $h_{2}$.
Table 5: Test accuracy (%) for different values of $\gamma$ on the USPS dataset with $h_{2}=10$ and $h_{1}=10^{-15}$
Method $\gamma=0.2$ $\gamma=0.6$ $\gamma=1.0$
GeoClip ($k=50$) 85.537 86.182 85.752
Table 6: Test accuracy (%) for different values of $h_{2}$ on the USPS dataset with $h_{1}=10^{-15}$ and $\gamma=1$
Method $h_{2}=1$ $h_{2}=10$ $h_{2}=100$ $h_{2}=1000$
GeoClip ($k=50$) 86.021 85.752 85.752 85.752
Table 7: Test accuracy (%) for different values of $h_{1}$ on the USPS dataset with $h_{2}=10$ and $\gamma=1$
Method $h_{1}=10^{-6}$ $h_{1}=10^{-10}$ $h_{1}=10^{-15}$
GeoClip ($k=50$) 85.752 85.752 85.752